OpenStudy (anonymous):

According to Car and Driver, an Alfa Romeo going at 70 mph requires 177 feet to stop. Assuming that the stopping distance is proportional to the square of velocity, find the stopping distances required by an Alfa Romeo going at 35 mph and at 140 mph (its top speed).

OpenStudy (anonymous):

proportional ratios: (V^2/distance) so, (70^2/177) = (V^2/35) isolate Velocity as: sqrt((70^2/177) * 35) = V and for the other speed of 140 mph replace 35 with 140 sqrt((70^2/177) * 140) = V